OpenStudy (freemap):

Which shows 3x + 6y = –12 in slope-intercept form? A. y=-2x-12 B. y=1/2x-12 C. y=-1/2x-2 2 D. y=2x-2

OpenStudy (freemap):

My answer is A

OpenStudy (freemap):

@radar

OpenStudy (anonymous):

not a

OpenStudy (anonymous):

3\6 = 1\2

OpenStudy (anonymous):

mine is c

OpenStudy (anonymous):

move 3x to other side divide by 6

OpenStudy (freemap):

That makes sense

OpenStudy (anonymous):

so it will be 6y+-3x -12

OpenStudy (anonymous):

6y=-3x-12

OpenStudy (anonymous):

y=3\6-12\6

OpenStudy (anonymous):

y=-1\2x-2

OpenStudy (freemap):

6=-12=-2 -3x=-12=4 2/4=1/2

OpenStudy (anonymous):

i just got done with slope intercept

OpenStudy (freemap):

finding slope you just substitute in 0 for one of variables and solve

OpenStudy (anonymous):

in my class lol

OpenStudy (radar):

Your Choice C has too many 2's in it. Check it for a typo.

OpenStudy (anonymous):

yes i know we also didthis and we had that question

OpenStudy (anonymous):

the right awnser is c tho

OpenStudy (lynfran):

slope intercept form y=mx+b 3x+6y=-12 6y=-3x-12 6y= -3(x+4) y=-3(x+4)/6 y=-1/2(x+4) y=-1/2x-2

OpenStudy (freemap):

The picture is what C is supposed to be

OpenStudy (anonymous):

lynfran thanks

OpenStudy (radar):

Congratulations you've done it.

OpenStudy (freemap):

Oh wow I didn't see what everybody typed, I was to busy typing myself. Thanks so so much for everybody who help me

OpenStudy (anonymous):

np

OpenStudy (radar):

LynFran has a step by step solution.

OpenStudy (anonymous):

yep i tryed to do that to but it didnt work out much lol

OpenStudy (freemap):

As I seen wonderful explanation, I knew how to find slope but not the other part

OpenStudy (freemap):

@blasterjakz you helped me to understand as well. I really appreciate it

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i need help but i cant get any from the english depo

OpenStudy (freemap):

I really wish I could help you there, but I don't want to give any wrong help and answers. I'm really sorry. Ask some of the like mentors, etc maybe they can help

OpenStudy (anonymous):

When The War of the Worlds was written, England feared _____. A. economic collapse B. plague C. alien invasion D. attack from another country

OpenStudy (anonymous):

need help with that

OpenStudy (freemap):

How about you ask @LynFran. Shes probably much better than I am. I'm not in your class, sorry

OpenStudy (anonymous):

@LynFran.

OpenStudy (anonymous):

can u hel p plz

OpenStudy (lynfran):

this is what google give The War of the Worlds thrilled and ... sensationalism, scientific fantasy, suburban mundanity and fears of invasion

OpenStudy (anonymous):

no one on english depo care atm